Problem 8-4
Expected and required rates of return

Assume that the risk-free rate is 3.5% and the market risk premium is 6%.

  1. What is the required return for the overall stock market? Round your answer to two decimal places.
    %

  2. What is the required rate of return on a stock with a beta of 2.4? Round your answer to two decimal places.
    %
Answer #1

Market risk premium=Market rate-Risk free rate

1.Market rate=(6+3.5)=9.5%

2.required return= risk-free rate +Beta*(market rate- risk-free rate )

=3.5+2.4*(9.5-3.5)=17.9%